OpenStudy (anonymous):

The fraction (1+(1/x))/(1-(1/x^2)) is equivalent to... please explain how you get your answer.

OpenStudy (anonymous):

\[\frac{ 1+\frac{ 1 }{ x } }{ 1-\frac{ 1 }{ x ^{2} } }\] thats the fraction if it makes it easier^

OpenStudy (anonymous):

are you supposed to get rid of the fraction?

OpenStudy (anonymous):

i mean turn it into a non fraction?

OpenStudy (anonymous):

no you are supposed to simplify the fraction

OpenStudy (anonymous):

ok so what is the least common denominator? for everything?

OpenStudy (anonymous):

x?

OpenStudy (anonymous):

close,

OpenStudy (anonymous):

still missing something

OpenStudy (anonymous):

\[x^{2}\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

so multiply each numerator and denominator by whats needed to get x^2 as the denominator for each term, 2 terms up top, 2 terms on bottom

OpenStudy (anonymous):

and you should get a fraction plus a fraction divided by a fraction minus a fraction

OpenStudy (anonymous):

each of the 4 fractions will have x^2 as the denominiator

OpenStudy (anonymous):

post that when you have it

OpenStudy (anonymous):

you know what i mean?

OpenStudy (anonymous):

ok guess not?

OpenStudy (anonymous):

im sorry i understand my computer just freaked out on me but i got it working again

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so\[\frac{ \frac{ x ^{2} }{ x ^{2} }+\frac{ x }{ x ^{2} } }{ \frac{ x ^{2} }{ x ^{2} }- \frac{ 1 }{ x ^{2} } }\]?

OpenStudy (anonymous):

thats correct, now combine the fractions in the numerator and combine the fractions in the denominator

OpenStudy (anonymous):

ok so...\[\frac{ \frac{ x ^{2}+x }{ x ^{2} } }{ \frac{ - x ^{2} }{ x ^{2} } }\]?

OpenStudy (anonymous):

the top looks right, but the bottom is (x^2 -1)/(x^2)

OpenStudy (anonymous):

after fixing that, invert and multiply the denominator and cancel the x^2's and then factor the new denominator, and cancel as necessary

OpenStudy (anonymous):

oh ok so...\[\frac{ \frac{ x(x+1) }{ x ^{2} } }{ \frac{ (x+1)(x-1) }{ x ^{2} } }\] then...\[\frac{ \frac{ x }{ x ^{2} } }{ \frac{ x-1 }{ x ^{2} } }\]? then what?

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

thats right but you can cancel out the x^2 leaving you with x/(x-1)

OpenStudy (anonymous):

remember though they are only equal when x != 0 and x != 1

OpenStudy (anonymous):

or is it...\[\frac{ x(x+1) }{ x ^{2} }\times \frac{ x ^{2} }{ (x+1)(x-1) }\] then you cross out (x+1)s and the x^2s and get... \[\frac{ x }{ (x+1) }\]

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

i meant \[\frac{ x }{ x-1 }\]

OpenStudy (anonymous):

yw :)